Jocelynnn:

Solve for x: 4x + 9 = 12.

xxDeppressionxx:

4x+9=12 equation solution: 12 = 4x + 9 12 - 9 = 4x 3x = 4 (divide both sides by 4 to get x) 3/4 x = (0.75) x = 4x/4